a new business on the first day he open his business made $520.00 on the fifth day he is $840.00 if his business countin at this same rate what will bring in on the tenth day
Answer by ewatrrr(24785)

a5 = 520 + 4d = 840, d = (840-520)/4 = $80
d = $80
a10 = 520 + 9($80) = $1240